OpenStudy (anonymous):

(x+3)(x+1)/(x+2)(x-3) partial fraction.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Apply the same method of the previuous problem

OpenStudy (anonymous):

complicated.. i dunno know what to do ??

OpenStudy (jojo4eva):

looks to me as if they are being divided is that correct?

OpenStudy (anonymous):

First do a division to find out that \[ \frac{(x+1) (x+3)}{(x-3) (x+2)}=\frac{5 x+9}{(x-3) (x+2)}+1 \] and after do the partial fraction for \[ \frac{5 x+9}{(x-3) (x+2)} \] like the one we did before.

OpenStudy (anonymous):

Your final answer should look like \[\frac{(x+1) (x+3)}{(x-3) (x+2)}=\frac{1}{5 (x+2)}+\frac{24}{5 (x-3)}+1 \]